# MaskedManipulator: Versatile Whole-Body Control for Loco-Manipulation

### v1

@@ Abstract.
Humans interact with their world while leveraging precise full-body control to achieve versatile goals. This versatility allows them to solve long-horizon, underspecified problems, such as placing a cup in a sink, by seamlessly sequencing actions like approaching the cup, grasping, transporting it, and finally placing it in the sink. Such goal-driven control can enable new procedural tools for animation systems, enabling users to define partial objectives while the system naturally “fills in” the intermediate motions. However, while current methods for whole-body dexterous manipulation in physics-based animation achieve success in specific interaction tasks, they typically employ control paradigms (e.g., detailed kinematic motion tracking, continuous object trajectory following, or direct VR teleoperation) that offer limited versatility for high-level goal specification across the entire coupled human-object system. To bridge this gap, we present MaskedManipulator, a unified and generative policy developed through a two-stage learning approach. First, our system trains a tracking controller to physically reconstruct complex human-object interactions from large-scale human mocap datasets. This tracking controller is then distilled into MaskedManipulator, which provides users with intuitive control over both the character’s body and the manipulated object. As a result, MaskedManipulator enables users to specify complex loco-manipulation tasks through intuitive high-level objectives (e.g., target object poses, key character stances), and MaskedManipulator then synthesizes the necessary full-body actions for a physically simulated humanoid to achieve these goals, paving the way for more interactive and life-like virtual characters.
copyright: none
@@ 1. Introduction
Generating versatile humanoid agents capable of autonomous and controllable behaviors is a shared challenge for both character animation and robotics. While recent approaches have significantly advanced locomotion capabilities for simulated (Tessler et al., 2024a; Wu et al., 2025) and real-world humanoids (Ji et al., 2024; Allshire et al., 2025; Ze et al., 2025), synthesizing controllers that can skillfully interact with the environment – especially with small, dynamic objects requiring precise dexterous manipulation – remains a daunting hurdle. A key challenge lies in how to effectively translate high-level human intent into low-level motor commands. For example, walking to an object, picking it up, inspecting the object, and then placing it back down on a table.
@@
We aim to build a versatile controller that seamlessly combines full-body locomotion with dexterous manipulation, which can also interpret such diverse command inputs. However, there exits an inherent tension: the need for broad task flexibility versus the need for exacting physical precision. On one hand, achieving versatility requires controllers capable of understanding abstract user intents, like sparsely defined target coordinates (e.g., “reach this point”) or high-level object goals (e.g., “place cup here”). These often underspecified goals present vast solution spaces, posing substantial challenges for control and learning techniques like reinforcement learning (RL). On the other hand, successful manipulation, particularly with common objects, demands meticulous precision. Even small deviations during physical interaction – grasping, repositioning, or placing – can lead to instability, object dropping, or complete task failure, often without recovery (Luo et al., 2024b). While prior work has tackled aspects of this challenge, existing methods typically excel at either sophisticated manipulation (Lin et al., 2024; Iyer et al., 2024) or versatile body control (e.g., via goal-conditioning (Tessler et al., 2024a)). Achieving unified versatile control that encompasses both body and object goals, and gracefully balances this flexibility-precision trade-off, remains an open problem.
@@
We propose a unified control framework for diverse and precise loco-manipulation tasks, responding to inputs ranging from detailed kinematic targets (e.g., tracking the hand positions via teleoperation) to sparse, high-level goals (like a desired position of an object). Our framework extends spatio-temporal goal-conditioning to encompass both the humanoid’s body parts and the manipulated object. To master the precise execution required by human-object interactions, our first stage (MimicManipulator) leverages human motion capture (e.g., GRAB dataset (Taheri et al., 2020)). The motion capture data demonstrates successful interaction strategies, such as sequences of grasping, placing, and regrasping objects, and hand-to-hand transfers. This physics-based tracker (MimicManipulator), trained with RL in a fully observable setting, learns the necessary actions to accurately reconstruct these intricate behaviors. Then, to enable versatile control from sparse goals, MimicManipulator’s expert interaction knowledge is distilled into our MaskedManipulator policy. By learning from MimicManipulator, MaskedManipulator becomes capable of producing human-like and diverse behaviors in response to under-specified goals for both body parts and objects (e.g., “bring the object to coordinate $(x,y,z)$”), effectively addressing the multi-solution challenge of versatile control.
@@
    A novel framework for versatile loco-manipulation that leverages human demonstrations to produce diverse, physically plausible, and human-like solutions to underspecified tasks involving coupled human-object interactions.
@@
    MimicManipulator: A physics-based motion tracker that accurately infers actions and reconstructs dexterous human loco-manipulation sequences from reference motion data.
@@
    MaskedManipulator: A unified generative loco-manipulation policy that enables diverse spatio-temporal goal-conditioning for both humanoid body parts and manipulated objects.
@@ 2. Related Work
Traditional reinforcement learning (RL) for character control often requires meticulous reward engineering. Imitation learning (IL) offers a powerful alternative by leveraging expert demonstrations. DeepMimic (Peng et al., 2018) showed RL can robustly replicate reference motions, a paradigm extended to object manipulation by works like (Yu et al., 2025) for dexterous sequences and InterMimic (Xu et al., 2025) for retargeting interactions across morphologies. Our MimicManipulator stage is inspired by these approaches, using RL to train a single policy to precisely track the entire diverse human loco-manipulation demonstrations from GRAB (Taheri et al., 2020), thereby recovering actions and ensuring physical plausibility.
@@
Integrating object manipulation significantly increases control complexity. Some approaches achieve manipulation without reference data via hand-crafted rewards and large-scale RL (Akkaya et al., 2019; Lin et al., 2024). More recently, learning-based methods using demonstrations have enabled full-body motion for object-trajectory tracking (Luo et al., 2024a). While these methods show progress in specific manipulation aspects, our work focuses on versatile, goal-conditioned manipulation from human data. MaskedManipulator uniquely allows explicit conditioning on the future state of the object itself as a goal, alongside humanoid body part goals. This enables synthesizing novel, full-body loco-manipulation behaviors from sparse constraints, offering a more flexible control paradigm.
@@
A trend in robotics and animation involves hierarchical control, where high-level systems specify abstract goals (e.g., walk targets (Peng et al., 2022), grasp poses (Li et al., 2024), 2D paths (Li et al., 2025)) for low-level execution, akin to System 1/System 2 reasoning (Kahneman, 2011). MaskedMimic (Tessler et al., 2024a) significantly advanced System1 versatility by framing character control as motion inpainting, allowing a unified model to synthesize diverse motions from sparse future joint targets. Our work extends this by enabling explicit spatio-temporal goal-conditioning on both humanoid body parts and the manipulated object within MaskedManipulator. This provides direct, versatile control over the entire coupled human-object system, crucial for sophisticated loco-manipulation tasks.
@@ 3. Preliminaries
We formulate our problem within the framework of goal-conditioned Markov Decision Process (MDP) (Puterman, 2014), defined by a tuple $M=(S,G,A,P,R,\gamma)$, consisting of states $S$, goals $G$, actions $A$, transition dynamics $P$, a reward function $R$, and a discount factor $\gamma$. The objective is to learn a policy $\pi(a_{t}|s_{t},g_{t})$ that selects an action $a_{t}$ based on the current state $s_{t}$ and a task-specific goal $g_{t}$.
The state of the humanoid’s $J$ links at time $t$ is described by their individual 3D positions $p_{j,t}\in\mathbb{R}^{3}$ orientations $\theta_{j,t}$, linear $v_{j,t}$ and angular $\omega_{j,t}$ velocities. Each orientation $\theta_{j,t}$ is represented using a continuous 6D rotation vector (Zhou et al., 2019). All link positions and orientations constitute the full body kinematic state, denoted $q_{t}=(\{p_{j,t}\}_{j=1}^{J},\{\theta_{j,t}\}_{j=1}^{J})$. The corresponding linear and angular velocities collectively form $\dot{q}_{t}$. When interacting with the scene, a body part may be subjected to contact forces, which we denote by $c_{j,t}$. Object states $q^{\text{obj}}_{t}$ are similarly defined by 3D position $p^{\text{obj}}_{t}$, 6D orientation $\theta^{\text{obj}}_{t}$, velocities $\dot{q}^{\text{obj}}_{t}$, and contact forces $c^{\text{obj}}_{t}$.
@@ 4. Method
Our approach for generating versatile humanoid loco-manipulation capabilities is structured around a two-stage learning process. Human motion capture data provides rich kinematic descriptions of complex interactions but typically lacks the underlying low-level motor commands (actions) required to drive a physics-based character. Our first stage, MimicManipulator, addresses these challenges by training a high-fidelity motion tracker within a physics-based simulation. Operating in a full-information setting with respect to the reference kinematics, MimicManipulator learns not only to recover the necessary actions but also to physically reconstruct the target loco-manipulation sequences, thereby ensuring dynamic feasibility and emphasizing the nuances of object handling. Subsequently, the robust interaction skills learned by MimicManipulator is distilled into a versatile control policy, MaskedManipulator. This second stage yields a policy that can generate novel, physically plausible behaviors by dynamically conditioning on spatio-temporal goals specified for various parts of the humanoid body or the manipulated object itself. The following sections detail the objectives, architecture, and training procedures for each of the two stages.
@@ 4.1. MimicManipulator
The first stage focuses on reproducing the intricate behaviors required for dexterous loco-manipulation. We train a single policy $\pi_{\text{track}}$, a motion-tracker denoted by MimicManipulator, to accurately reproduce the complex loco-manipulation sequences from the processed GRAB dataset. The reference kinematic data provides a rich description of the desired high-level behavior but does not contain the low-level motor commands (actions $a_{t}$) necessary to execute these motions in a physically simulated environment. Therefore, we formulate this task as a reinforcement learning problem where $\pi_{\text{track}}$ learns to infer these actions by imitating the reference motion.
@@ 4.1.1. Task design
To enable object interactions, the observation $s_{t}$ comprises of the character’s proprioception $q_{t}$, object information $q^{\text{obj}}_{t}$ and character-object relational features. All components are expressed in a character’s local coordinate frame and aligned with it’s current root heading (yaw) for policy invariance. The character-object relational features capture the spatial relationship between the humanoid and the object, by providing the vector (direction and distance) from each body part to the object surface. Object structure: Precise interaction requires awareness of the object geometry. To accomodate this, we also provide a Basis-Point-Set (Prokudin et al., 2019, BPS) representation, which represents all objects by the surface distances to a pre-sampled set of points. Compared to point-cloud representations, the pre-sampled and fixed BPS allow processing using simpler architectures (e.g., MLP), allowing faster training and inference, while still being expressive. Future reference trajectory (goal): To reproduce the target motions, MimicManipulator observes the the goal future poses in the next frame $K=1$ and $K=30$ frames (1 second), denoted by $g_{t}^{\text{track}}$. Each target future pose observation $\{\hat{q}_{t+k}\Theta q_{t}\}=\{\hat{p}_{t+k}-p_{t},\hat{q}_{t+k}\Theta q_{t}\}$ is a future pose $\hat{q}_{t+k}$ normalized with respect to the current simulated pose $q_{t}$, where $\Theta$ represents the inverse quaternion. In addition, we provide $\hat{c}_{t+k}$ to indicate which body parts should have contact with the object.
@@
The reward $R_{\text{track}}$ is a product of several components, ensuring high reward only when all aspects of the reference are accurately matched (Park et al., 2019; Won et al., 2020; Xu et al., 2025):
@@
Each component $r^{(\cdot)}$ is an exponential of a negatively weighted error or cost (e.g., $r=\exp(-w\cdot\text{cost})$), valued between (0, 1]. $r^{\text{humanoid}}=r^{\text{ht}}\cdot r^{\text{hr}}$ and $r^{\text{obj}}=r^{\text{ot}}\cdot r^{\text{or}}$ components consider the translation and rotation errors for the human and object respectively. The energy component attempts to mitigate artifacts such as jitters. The interaction component aims to prevent human-object penetrations by encouraging a gentle approach and interaction.
The contact component, Figure 2, encourages the policy to replicate the reference contact patterns and timing throughout the interaction sequence. It guides the agent through: Approach & engagement: Smoothly guiding the end-effectors (e.g., hands) towards the target contact points on the object, matching the reference approach. Maintaining contact: Ensuring stable contact once established, consistent with the demonstration. Disengagement: Facilitating a clean and timely release, mirroring the reference disengagement.
We found that the multiplicative design helps learning precise, long-horizon loco-manipulation by strictly requiring competence in all elements of the complex motion. Complementing the reward design, we leverage early termination to define a feasibility envelope for the training phase (Peng et al., 2018; Luo et al., 2023; Tessler et al., 2024a). An episode is terminated if any body part deviates by ¿25cm from reference, or the object deviates ¿10cm from reference, or when there’s a prolonged unintended loss of contact for ¿10 consecutive frames during a contact phase, or if the character maintains contact ¿0.4sec after the reference has released.
@@
We train a single policy to reconstruct the entire dataset. The reference sequences span a wide range of difficulty. This variation arises from both the object’s characteristics (e.g., grasping a mug by its handle is more demanding than grasping a banana) and the intricacy of the interaction itself (e.g., bimanual operations or tasks requiring precise tool-like dexterity). During training we periodically evaluate the agent’s performance and over-sample sequences in which it has failed (Luo et al., 2024a; Tessler et al., 2024a; Luo et al., 2023).

### v3

# MaskedManipulator: Versatile Whole-Body Manipulation

@@ Abstract.
We tackle the challenges of synthesizing versatile, physically simulated human motions for full-body object manipulation. Unlike prior methods that are focused on detailed motion tracking, trajectory following, or teleoperation, our framework enables users to specify versatile high-level objectives such as target object poses or body poses. To achieve this, we introduce MaskedManipulator, a generative control policy distilled from a tracking controller trained on large-scale human motion capture data. This two-stage learning process allows the system to perform complex interaction behaviors, while providing intuitive user control over both character and object motions. MaskedManipulator produces goal-directed manipulation behaviors that expand the scope of interactive animation systems beyond task-specific solutions.
submissionid: 126journalyear: 2025copyright: ccconference: SIGGRAPH Asia 2025 Conference Papers; December 15–18, 2025; Hong Kong, Hong Kongbooktitle: SIGGRAPH Asia 2025 Conference Papers (SA Conference Papers ’25), December 15–18, 2025, Hong Kong, Hong Kongdoi: 10.1145/3757377.3763934isbn: 979-8-4007-2137-3/2025/12ccs: Computing methodologies Physical simulationccs: Computing methodologies Control methodsccs: Computing methodologies Reinforcement learning
@@ 1. Introduction
Generating versatile humanoid agents capable of autonomous and controllable behaviors is a shared challenge for both character animation and robotics. While recent approaches have significantly advanced locomotion capabilities for simulated (tessler2024maskedmimic; wu2025uniphys) and real-world humanoids (ji2024exbody2; allshire2025visual; ze2025twist), synthesizing controllers that can skillfully interact with the environment – especially with small, dynamic objects requiring precise dexterous manipulation – remains a daunting hurdle. A key challenge lies in how to effectively translate high-level human intent into low-level motor commands. For example, walking to an object, picking it up, inspecting the object, and then placing it back down on a table.
@@
We aim to build a versatile controller that seamlessly combines full-body locomotion with dexterous manipulation, which can also interpret such diverse command inputs. However, there exits an inherent tension: the need for broad task flexibility versus the need for exacting physical precision. On one hand, achieving versatility requires controllers capable of understanding abstract user intents, like sparsely defined target coordinates (e.g., “reach this point”) or high-level object goals (e.g., “place cup here”). These often underspecified goals present vast solution spaces, posing substantial challenges for control and learning techniques like reinforcement learning (RL). On the other hand, successful manipulation, particularly with common objects, demands meticulous precision. Even small deviations during physical interaction – grasping, repositioning, or placing – can lead to instability, object dropping, or complete task failure, often without recovery (luo2024precise). While prior work has tackled aspects of this challenge, existing methods typically excel at either sophisticated manipulation (lin2024twisting; iyer2024open) or versatile body control (e.g., via goal-conditioning (tessler2024maskedmimic)). Achieving unified versatile control that encompasses both body and object goals, and gracefully balances this flexibility-precision trade-off, remains an open problem.
@@
We propose a unified control framework for diverse and precise full-body-manipulation tasks, responding to inputs ranging from detailed kinematic targets (e.g., tracking the hand positions via teleoperation) to sparse, high-level goals (like a desired position of an object). Our framework extends spatio-temporal goal-conditioning (MaskedMimic (tessler2024maskedmimic)) to encompass both the humanoid’s body parts and the manipulated object. To master the precise execution required by human-object interactions, our first stage (MimicManipulator) leverages human motion capture (e.g., GRAB dataset (taheri2020grab)). The motion capture data demonstrates successful interaction strategies, such as sequences of grasping, placing, and regrasping objects, and hand-to-hand transfers. This physics-based tracker (MimicManipulator), trained with RL in a fully observable setting, learns the necessary actions to accurately reconstruct these intricate behaviors. Then, to enable versatile control from sparse goals, MimicManipulator’s expert interaction knowledge is distilled into our MaskedManipulator policy. By learning from MimicManipulator, MaskedManipulator becomes capable of producing human-like and diverse behaviors in response to under-specified goals for both body parts and objects (e.g., “bring the object to coordinate (x,y,z)(x,y,z)”), effectively addressing the multi-solution challenge of versatile control.
@@
    We extend the versatile full-body controller, MaskedMimic (tessler2024maskedmimic), enabling full-body-manipulation. We demonstrate how to successfully leverage human demonstrations to produce diverse, physically plausible, and human-like solutions to underspecified tasks involving coupled human-object interactions.
@@
    MimicManipulator: A physics-based motion tracker that accurately infers actions and reconstructs dexterous human full-body-manipulation sequences from reference motion data.
@@
    MaskedManipulator: A unified generative full-body-manipulation policy that enables diverse spatio-temporal goal-conditioning for both humanoid body parts and manipulated objects.
@@ 2. Related Work
Traditional reinforcement learning (RL) for character control often requires meticulous reward engineering. Imitation learning (IL) offers a powerful alternative by leveraging expert demonstrations. DeepMimic (peng2018deepmimic) showed RL can robustly replicate reference motions, a paradigm extended to object manipulation by works like (yu2025skillmimic) for dexterous sequences and InterMimic (xu2025intermimic) for retargeting interactions across morphologies. Our MimicManipulator stage is inspired by these approaches, using RL to train a single policy to precisely track the entire diverse human full-body-manipulation demonstrations from GRAB (taheri2020grab), thereby recovering actions and ensuring physical plausibility.
@@
Integrating object manipulation into humanoid control substantially increases task complexity. Early approaches achieved progress without reference data by relying on hand-crafted rewards and large-scale reinforcement learning (akkaya2019solving; lin2024twisting). More recent work has explored demonstration-driven methods: for example, OmniGrasp (luo2024omnigrasp) synthesizes full-body motions conditioned on dense object trajectories and trained via dense rewards. It achieves this by leveraging a scene-agnostic generative prior that is steered through hierarchical reinforcement learning. Compared to our approach, this design presents two key limitations: (1) the reliance on dense reward signals restricts OmniGrasp to short-horizon training objectives, making it unsuitable for more general goal-conditioned tasks such as bringing an object to a target position (pignatelli2023survey); (2) the scene-agnostic prior hinders dexterous control. As a result, OmniGrasp struggles reproducing dexterous behaviors such as grasping a teapot by its handle or performing hand-to-hand transfers. In contrast, our method first learns to reproduce dexterous human-object demonstrations from dense goals and then distills this knowledge into a flexible, long-horizon, goal-conditioned controller. The resulting controller, MaskedManipulator, supports a range of object and/or humanoid body-part goals, enabling the synthesis of diverse full-body-manipulation behaviors from sparse constraints and supporting more adaptable, long temporal sequence control.
@@
A trend in robotics and animation involves hierarchical control, where high-level systems specify abstract goals (e.g., walk targets (peng2022ase), grasp poses (li2024virt), 2D paths (li2025hamster)) for low-level execution, akin to System 1/System 2 reasoning (kahneman2011thinking). MaskedMimic (tessler2024maskedmimic) significantly advanced System1 versatility by framing character control as motion inpainting, allowing a unified model to synthesize diverse motions from sparse future joint targets. Our work extends this by enabling explicit spatio-temporal goal-conditioning on both humanoid body parts and the manipulated object within MaskedManipulator. This provides direct, versatile control over the entire coupled human-object system, crucial for sophisticated full-body-manipulation tasks.
@@ 3. Preliminaries
We formulate our problem within the framework of goal-conditioned Markov Decision Process (MDP) (puterman2014markov), defined by a tuple $M=(S,G,A,P,R,\gamma)$, consisting of states $S$, goals $G$, actions $A$, transition dynamics $P$, a reward function $R$, and a discount factor $\gamma$. The objective is to learn a policy $\pi(a_{t}|s_{t},g_{t})$ that selects an action $a_{t}$ based on the current state $s_{t}$ and a task-specific goal $g_{t}$.
The state of the humanoid’s $J$ links at time $t$ is described by their individual 3D positions $p_{j,t}\in\mathbb{R}^{3}$ orientations $\theta_{j,t}$, linear $v_{j,t}$ and angular $\omega_{j,t}$ velocities. Each orientation $\theta_{j,t}$ is represented using a continuous 6D rotation vector (zhou2019continuity). All link positions and orientations constitute the full body kinematic state, denoted $q_{t}=(\{p_{j,t}\}_{j=1}^{J},\{\theta_{j,t}\}_{j=1}^{J})$. The corresponding linear and angular velocities collectively form $\dot{q}_{t}$. When interacting with the scene, a body part may be subjected to contact forces, which we denote by $c_{j,t}$. Object states $q^{\text{obj}}_{t}$ are similarly defined by 3D position $p^{\text{obj}}_{t}$, 6D orientation $\theta^{\text{obj}}_{t}$, velocities $\dot{q}^{\text{obj}}_{t}$, and contact forces $c^{\text{obj}}_{t}$.
@@ 4. Method
Our approach for generating versatile humanoid full-body-manipulation capabilities is structured around a two-stage learning process. Human motion capture data provides rich kinematic descriptions of complex interactions but typically lacks the underlying low-level motor commands (actions) required to drive a physics-based character. Our first stage, MimicManipulator, addresses these challenges by training a high-fidelity motion tracker within a physics-based simulation. Operating in a full-information setting with respect to the reference kinematics, MimicManipulator learns not only to recover the necessary actions but also to physically reconstruct the target full-body-manipulation sequences, thereby ensuring dynamic feasibility and emphasizing the nuances of object handling. Subsequently, the robust interaction skills learned by MimicManipulator is distilled into a versatile control policy, MaskedManipulator. This second stage yields a policy that can generate novel, physically plausible behaviors by dynamically conditioning on spatio-temporal goals specified for various parts of the humanoid body or the manipulated object itself. The following sections detail the objectives, architecture, and training procedures for each of the two stages.
@@ 4.1. MimicManipulator
The first stage focuses on reproducing the intricate behaviors required for dexterous full-body-manipulation. We train a single policy $\pi_{\text{track}}$, a motion-tracker denoted by MimicManipulator, to accurately reproduce the complex full-body-manipulation sequences from the processed GRAB dataset. The reference kinematic data provides a rich description of the desired high-level behavior but does not contain the low-level motor commands (actions $a_{t}$) necessary to execute these motions in a physically simulated environment. Therefore, we formulate this task as a reinforcement learning problem where $\pi_{\text{track}}$ learns to infer these actions by imitating the reference motion.
@@ 4.1.1. Task design
To enable object interactions, the observation $s_{t}$ comprises of the character’s proprioception $q_{t}$, object information $q^{\text{obj}}_{t}$ and character-object relational features. All components are expressed in a character’s local coordinate frame and aligned with it’s current root heading (yaw) for policy invariance. The character-object relational features capture the spatial relationship between the humanoid and the object, by providing the vector (direction and distance) from each body part to the object surface. Object structure: Precise interaction requires awareness of the object geometry. To accomodate this, we also provide a Basis-Point-Set (prokudin2019efficient, BPS) representation, which represents all objects by the surface distances to a pre-sampled set of points. Compared to point-cloud representations, the pre-sampled and fixed BPS allow processing using simpler architectures (e.g., MLP), allowing faster training and inference, while still being expressive. Future reference trajectory (goal): To reproduce the target motions, MimicManipulator observes the the goal future poses in the next frame $K=1$ and $K=30$ frames (1 second), denoted by $g_{t}^{\text{track}}$. Each target future pose observation $\{\hat{q}_{t+k}\Theta q_{t}\}=\{\hat{p}_{t+k}-p_{t},\hat{q}_{t+k}\Theta q_{t}\}$ is a future pose $\hat{q}_{t+k}$ normalized with respect to the current simulated pose $q_{t}$, where $\Theta$ represents the inverse quaternion. In addition, we provide $\hat{c}_{t+k}$ to indicate which body parts should have contact with the object.
@@
The reward $R_{\text{track}}$ is a product of several components. We found that the multiplicative design helps learning precise, full-body-manipulation over long temporal sequences by strictly requiring competence in all elements of the complex motion (park2019learning; won2020scalable; xu2025intermimic).
@@
Each component $r^{(\cdot)}$ is an exponential of a negatively weighted error or cost (e.g., $r=\exp(-w\cdot\text{cost})$), valued between (0, 1].
(2) $r^{\text{humanoid}}=r^{\text{ht}}\cdot r^{\text{hr}}\,\,\text{and}\,\,r^{\text{obj}}=r^{\text{ot}}\cdot r^{\text{or}}$
components consider the translation and rotation errors for the human and object respectively. The energy component attempts to mitigate artifacts such as jitters. The interaction component aims to prevent human-object penetrations by encouraging a gentle approach and interaction.
The contact component, illustrated in Figure 2, is formulated to encourage the policy to reproduce the contact patterns and timing observed in the reference motion. The reward is structured into three distinct phases, each corresponding to a specific stage of the interaction, thereby guiding the agent to accurately replicate the demonstrated behavior. To prevent abrupt behaviors, the approach and engagement phase starts 1 second before contact and ends 0.2 seconds after the reference motion indicates contact has started. Similarly, the disengagement phase beins 1 second before the reference motion indicates that contact has ended. We provide additional details in the supplementary material.
In the approach and engagement phase, we identify future contact bodies from the reference trajectory. For each future contact body, the closest point on the object mesh is determined in both the simulation and the reference motion. The reward then penalizes discrepancies in distances and surface normals between each simulated and reference contact body, thereby promoting similarity in the approach trajectory and orientation. This mechanism ensures that end-effectors (e.g., hands) are guided smoothly toward the target contact regions in a manner consistent with the demonstration.
During the contact maintenance phase, we identify which reference body parts $\hat{c}_{j,t}$ are marked as should-be-in-contact. For each such body part, a reward is computed based on its proximity to the object mesh. This formulation encourages stable contact, discourages premature detachment, and ensures alignment of the contact configuration with that of the reference motion.
Finally, in the disengagement phase, previously contacting body parts are incentivized to move away from the object in a controlled and timely manner. Rewards are again based on distance and surface normal alignment, encouraging a smooth release that mirrors the reference disengagement dynamics.
Complementing the reward design, we leverage early termination to define a feasibility envelope for the training phase (peng2018deepmimic; luo2023perpetual; tessler2024maskedmimic). An episode is terminated if any body part deviates by ¿25cm from reference, or the object deviates ¿10cm from reference, or when there’s a prolonged unintended loss of contact for ¿10 consecutive frames during a contact phase, or if the character maintains contact ¿0.4sec after the reference has released.
@@
We train a single policy to reconstruct the entire dataset. The reference sequences span a wide range of difficulty. This variation arises from both the object’s characteristics (e.g., grasping a mug by its handle is more demanding than grasping a banana) and the intricacy of the interaction itself (e.g., bimanual operations or tasks requiring precise tool-like dexterity). During training we periodically evaluate the agent’s performance and over-sample sequences in which it has failed (luo2024omnigrasp; tessler2024maskedmimic; luo2023perpetual).